OpenStudy (anonymous):

If a spherical ball is enlarged so that its surface area is 9 times greater than its original surface area, then the original radius was multiplied by _________. Answer

OpenStudy (campbell_st):

try a simple experiment r = 1 \[SA = 4\pi \times 1^2\] \[SA = 4 \pi\] then 9 times the SA \[36\pi = 4\pi \times r^2\] \[r^2 = 9\] the radius is 3 so the length of the radius is tripled

OpenStudy (anonymous):

ooh i see so wait that was just an example right?

OpenStudy (campbell_st):

no its the solution but for area the linear ratio is squared radius 1:3 for area becomes 1^2 : 3^2 or 1: 9 volume is linear ratio cubed

OpenStudy (campbell_st):

I gave you a simple way of solving the problem by letting the radius =1 then multiplying the surface area by 9 and then solving for r

OpenStudy (anonymous):

oh ok. thank-you